OpenStudy (anonymous):

what are the zeros in this functions y=x squared-x-20 show work!

OpenStudy (amistre64):

5 and -4 :)

OpenStudy (amistre64):

(a+b)(a+c) = aa + ac + ba + bc a^2 +(b+c)a + (bc)

OpenStudy (amistre64):

-5+4 = -1 -5(4) = -20

OpenStudy (anonymous):

thank you! u are big help